My mum is 3 times as old as I am. If she is 24 years older than me, what are our ages?
Answer by stanbon(75887)

My mum is 3 times as old as I am.
If she is 24 years older than me, what are our ages?
------------
Let mum = m
Let self = s
----
Equations:
m = 3s
m = s + 24
-----
3s = s+24
2s = 24
s = 12 (Self's age now)
----
m = 3s = 36 (mum's age now)
===============================
Cheers,
Stan H.
=================